OpenStudy (anonymous):

x^2 - 5x - 14 = 0 Can someone please explain this. I have to use the quadratic formula.

OpenStudy (mertsj):

a=1, b=-5, c=-14

OpenStudy (mertsj):

Look up the quadratic formula and replace each a with 1, each b with -5, and each c with -14. Simplify.

OpenStudy (anonymous):

\[\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] \[ax^2 + bx + c = 0\]

OpenStudy (anonymous):

you got this?

OpenStudy (anonymous):

course it would be a hell of a lot easier if you wrote \[(x-7)(x+2)=0\]and solved that one

OpenStudy (mertsj):

Unless, of course, the directions say to use the formula.

OpenStudy (anonymous):

i think it should say "complete the square"

OpenStudy (anonymous):

:/ it doesn't It just said i had to use the quad. formula.

OpenStudy (mertsj):

Or "solve by factoring".

OpenStudy (mertsj):

That is designed to give you practice with the formula on the simple problems.

OpenStudy (anonymous):

\[ax^2+bx+c=0\\x^2-5x-14=0\] so \(a=1, b=-5,c=-14\) substitute in to \[\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] and get \[\frac{-(-5) \pm \sqrt{(-5)^2-4\times 1\times (-14)}}{2\times 1}\] then do careful arithmetic

OpenStudy (anonymous):

thank you i appreciate this

OpenStudy (anonymous):

you will know you are right when you get \(-2\) or \(7\) if you get stuck let us know

OpenStudy (anonymous):

Im good from here:-)

OpenStudy (anonymous):

k